# On local Galois deformation rings: generalised reductive groups

### v1

@@ 4.1. Continuity via algebra
We first focus on the case, when $f$ is a monomial of degree $d\geq 1$. There exists an $A$-multilinear form $\beta:A^{\otimes_{A}d}\to A$, such that $f(a)=\beta(a\otimes\dots\otimes a)$ for all $a\in A$. Write $\iota:M\to A$ for the inclusion map. We have a natural $R$-linear map $\iota^{\otimes}:M^{\otimes_{R}d}\to A^{\otimes_{A}d}$ and we denote by $N$ the image of $M^{\otimes_{R}d}$ under $\tilde{\beta}:=\beta\circ\iota^{\otimes}$ in $A$. In particular $f|_{M}$ takes values in $N$ and arises as the composition
@@
Let $X$ be an affine scheme of finite type over $\mathcal{O}$. A point $x\in X(A)$ corresponds to an $\mathcal{O}$-algebra homomorphism $x:B\rightarrow A$. We thus have a canonical map
@@
    for all $n\geq 1$ and all closed immersions $\tau:X\hookrightarrow\mathbb{A}^{n}$ of $\mathcal{O}$-schemes there exists a finitely generated $R$-submodule $M\subseteq\mathbb{A}^{n}(A)=A^{n}$ such that $\tau(f(S))\subseteq M$ and $\tau\circ f:S\rightarrow M$ is continuous for a profinite topology on the target.
@@
    there exists a closed immersion $\tau:X\hookrightarrow\mathbb{A}^{n}$ of $\mathcal{O}$-schemes and a finitely generated $R$-submodule $M\subseteq\mathbb{A}^{n}(A)=A^{n}$ such that $\tau(f(S))\subseteq M$ and $\tau\circ f:S\rightarrow M$ is continuous for a profinite topology on the target.
@@
[…]
@@
Let us assume that (3) holds. Let $M_{i}\subseteq A$ be the image of $M\subseteq\prod_{i=1}^{n}A$ under the projection onto the $i$-th component. Then $M\subseteq\prod_{i=1}^{n}M_{i}$ and $\prod_{i=1}^{n}M_{i}$ is a finitely generated $R$-module. Hence, we may replace $M$ with $\prod_{i=1}^{n}M_{i}$. Let $b_{i}\in B$ be the images of $x_{i}$ as before. If $b\in B$ then there is $g\in\mathcal{O}[x_{1},\ldots,x_{n}]$ such that $b=g(b_{1},\ldots,b_{n})$. The image of $\prod_{i=1}^{n}M_{i}$ under the map $\varphi_{g}:A^{n}\rightarrow A$, $(a_{1},\ldots,a_{n})\mapsto g(a_{1},\ldots,a_{n})$ is contained in a finitely generated $R$-module $M_{b}$ by the arguments in Lemma 4.1. Since $\tau(f(S))\subseteq\prod_{i=1}^{n}M_{i}$ we have $\varphi_{g}(\tau(f(S)))\subseteq M_{b}$. Moreover, the induced map $S\rightarrow M_{b}$ is continuous since polynomial maps are continuous. The composition $\varphi_{g}\circ\tau:X(A)\rightarrow A$ is the map given by evaluation at $b$. Hence, (3) implies (1). ∎
@@ 4.2. Continuity via condensed sets
We will now express the equivalent conditions in Lemma 4.2 in terms of condensed sets. Recall that these are certain functors from the category of profinite sets to the category of sets, see [48] and Appendix A. We may view profinite sets as condensed sets via the Yoneda embedding $S\mapsto\underline{S}$, where $\underline{S}(T):=\mathcal{C}(T,S)$. The underlying set of a condensed set is just that functor evaluated at a point.
Recall further, that every profinite group (ring) is naturally a condensed group (ring), as the functor from topological spaces to condensed sets commutes with finite limits [48, Proposition 1.7]. We can also equip a profinite ring $R$ with the discrete topology and write $R_{\operatorname{disc}}$ to denote the associated discrete condensed ring. We have a natural map of condensed rings $R_{\operatorname{disc}}\to\underline{R}$. An abstract $R$-module $M$ can be seen as a discrete condensed $R_{\operatorname{disc}}$-module $M_{\operatorname{disc}}$, so that $M_{\operatorname{disc}}(T)=\mathcal{C}(T,M)$, where the target $M$ is given the discrete topology. The tensor product of condensed $R_{\operatorname{disc}}$-modules $M_{\operatorname{disc}}\otimes_{R_{\operatorname{disc}}}\underline{R}$ is a condensed $R$-module. [^2] Fargues–Scholze in [28, Section VIII] call such modules relatively discrete $R$-modules.
@@
We will say that a map of sets $f:S\rightarrow X(A)$ is $R$-condensed if there is morphism of condensed sets $\tilde{f}:\underline{S}\to X(A_{\operatorname{disc}}\otimes_{R_{\operatorname{disc}}}\underline{R})$, whose map on underlying sets is $f$.
@@
By A.8 the map $X(A_{\operatorname{disc}}\otimes_{R_{\operatorname{disc}}}\underline{R})\to\mathbb{A}^{n}(A_{\operatorname{disc}}\otimes_{R_{\operatorname{disc}}}\underline{R})$ is a quasi-compact injection. It follows from [47, Proposition 4.13], that
@@ 4.3. $R$-condensed representations
[…]
@@
(1) implies (2). Let $\gamma_{1},\dots,\gamma_{N}\in\Gamma$ be generators of a dense subgroup of $\Gamma$. By assumption there is an $R$-subalgebra $B\subseteq A$, which contains an $R$-finite open $R$-subalgebra $B^{0}$, such that $\rho(\gamma_{1}),\dots,\rho(\gamma_{N})\in G(B)$. Since $G(B)$ is closed in $G(A)$, we have $\rho(\Gamma)\subseteq G(B)$. Since $G$ is of finite type over $\mathcal{O}$, $G(B^{0})$ is an open subgroup of $G(B)$. Thus the preimage $\Gamma^{\prime}:=\rho^{-1}(G(B_{0}))$ is an open subgroup of $\Gamma$. If $\tau:G\rightarrow\mathbb{A}^{n}$ is a closed immersion then $\tau(\rho(\Gamma^{\prime}))$ is contained in $\mathbb{A}^{n}(B^{0})$, which a finitely generated $R$-module. Thus the restriction of $\rho$ to $\Gamma^{\prime}$ is $R$-condensed and Lemma 4.13 implies that $\rho$ is $R$-condensed.
@@
Clearly (1) implies (2). For the converse, let $\Delta$ be an abstract subgroup of $\Gamma$ generated by $\{\gamma_{i}\}_{i\in I}$. Let $\tau:G\hookrightarrow\mathbb{A}^{n}$ be a closed immersion. Since $\rho$ is $R$-condensed $\rho(\Gamma)$ is contained in a finitely generated $R$-submodule $M\subseteq\mathbb{A}^{n}(A)$ and the map $\rho:\Gamma\rightarrow M$ is continuous. The assumption in (2) implies that $\rho(\Delta)\subseteq H(A)$. Since $H$ is closed in $G$ the restriction of $\tau$ to $H$ is a closed immersion. It follows from Lemma 4.1 that $M\cap H(A)$ is a closed subset of $M$. Its preimage in $\Gamma$ will be a closed subset of $\Gamma$ containing $\Delta$. Since $\Delta$ is dense in $\Gamma$ we conclude that $\rho(\Gamma)\subseteq M\cap H(A)$, which implies part (1). ∎
@@ 5. The space of generic matrices
Let $G$ be a smooth affine $\mathcal{O}$-group scheme and let $\overline{\rho}:\Gamma\rightarrow G(k)$ be a continuous representation, where $\Gamma$ is a profinite group satisfying Mazur’s $p$-finiteness condition. We will introduce a space of generic matrices for $G$, which will depend on $\overline{\rho}$ and a faithful representation $\tau:G\to\operatorname{GL}_{d}$.
@@ 5.1. Definition of $X^{\mathrm{gen}}_{\operatorname{GL}_{d}}$ and $X^{\mathrm{gen},\tau}_{G}$
We fix [^3] Its existence follows from [30, Exp. $\mathrm{VII}_{B}$, Rem. 11.11.1]. a closed immersion $\tau:G\to\operatorname{GL}_{d}$ of $\mathcal{O}$-group schemes. By composing it with $\overline{\rho}$ we obtain a continuous representation $\tau\circ\overline{\rho}:\Gamma\rightarrow\operatorname{GL}_{d}(k)$. The linearisation $(\tau\circ\overline{\rho})^{\mathrm{lin}}:k[\Gamma]\to k$ extends to the completed group ring $(\tau\circ\overline{\rho})^{\mathrm{lin}}:k\llbracket\Gamma\rrbracket\to k$. Let $\overline{D}:k\llbracket\Gamma\rrbracket\rightarrow k$ be the continuous determinant law associated to $(\tau\circ\overline{\rho})^{\mathrm{lin}}$. (See [16] and [57, Section 3].)
Let $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$ be the universal deformation ring of $\overline{D}$ introduced in [16, Section 3.1] and let $D^{u}:R^{\mathrm{ps}}_{\operatorname{GL}_{d}}\llbracket\Gamma\rrbracket\rightarrow R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$ be the extension of the universal $d$-dimensional determinant law lifting $\overline{D}$. In particular, $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$ is a complete local noetherian $\mathcal{O}$-algebra with residue field $k$, and we equip it with its $\mathfrak{m}$-adic topology. We will write $D^{u}_{|A}$ for the specialisation of $D^{u}$ along $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}\rightarrow A$. We let $X^{\mathrm{ps}}_{\operatorname{GL}_{d}}:=\operatorname{Spec}(R^{\mathrm{ps}}_{\operatorname{GL}_{d}})$. Note, that $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$ corresponds to $R^{\mathrm{ps}}$ of [8] when we take the residual representation to be $\tau\circ\overline{\rho}$.
Let $\mathrm{CH}(D^{u})$ be the two-sided ideal of $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}\llbracket\Gamma\rrbracket$ defined in [16, Section 1.17], so that $E^{u}:=R^{\mathrm{ps}}_{\operatorname{GL}_{d}}\llbracket\Gamma\rrbracket/\mathrm{CH}(D^{u})$ is the largest quotient of $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}\llbracket\Gamma\rrbracket$ on which $D^{u}$ descends to a Cayley–Hamilton determinant law. Since $\Gamma$ satisfies Mazur’s $p$-finiteness condition, $E^{u}$ is a finitely generated $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$-module and $\mathrm{CH}(D^{u})$ is closed by [57, Proposition 3.6]. We write $D^{u}_{E^{u}}:E^{u}\to R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$ for the determinant law induced on $E^{u}$ by $D^{u}$ and $\pi_{E^{u}}:R^{\mathrm{ps}}_{\operatorname{GL}_{d}}\llbracket\Gamma\rrbracket\to E^{u}$ for the projection.
@@
It follows from [57, Proposition 3.6] that $E^{u}$ is a finitely generated $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$-module. This implies that the quotient topology on $E^{u}$ coincides with the topology inherited from $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$ as a finitely generated $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$-module, see the proof of [8, Lemma 3.2]. Every $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$-submodule of $E^{u}$ is closed with respect to this topology. Hence, $S$ is closed in $E^{u}$, and thus $S=E^{u}$. ∎
@@
(1) implies (2). The homomorphism $\widehat{\rho}\otimes A:R^{\mathrm{ps}}_{\operatorname{GL}_{d}}\llbracket\Gamma\rrbracket\otimes_{R^{\mathrm{ps}}_{\operatorname{GL}_{d}}}A\to M_{d}(A)$ induces a $d$-dimensional $A$-valued Cayley–Hamilton determinant law
@@
Part (1). Recall, that in [8, Lemma 3.1] the $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$-algebra $A^{\mathrm{gen}}$ is proved to represent the functor, that maps an $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$-algebra $A$ to the set of homomorphisms of Cayley–Hamilton algebras $f:E^{u}\to M_{d}(A)$. Once we prove this functor to be naturally isomorphic to $X^{\mathrm{gen}}_{\operatorname{GL}_{d}}$ as in 5.3, the claim about representability follows and $X^{\mathrm{gen}}_{\operatorname{GL}_{d}}$ is naturally isomorphic to the $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$-scheme $X^{\mathrm{gen}}=\operatorname{Spec}A^{\mathrm{gen}}$ from [8].
[…]
@@
Part (2) follows from [8, Lemma 3.2] or Lemma 4.9. ∎
@@
As already explained above, [57, Proposition 3.6] implies that $E^{u}$ is a finitely generated $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$-module, and hence $k\otimes_{R^{\mathrm{ps}}_{\operatorname{GL}_{d}}}E^{u}$ is a finite dimensional $k$-vector space.
@@
It follows from Part (1) of Lemma 5.4 that $X^{\mathrm{gen}}_{\operatorname{GL}_{d}}$ coincides with the scheme $X^{\mathrm{gen}}$, that was contructed in [8, Lemma 3.1] as a closed subscheme of $M_{d}^{n}\times X^{\mathrm{ps}}_{\operatorname{GL}_{d}}$ by choosing generators $\delta_{1},\ldots,\delta_{n}$ of $E^{u}$ as an $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$-module and mapping $\rho$ to the $n$-tuple $(\rho(\delta_{1}),\ldots,\rho(\delta_{n}))$. Since $\rho(\gamma_{i})$ are invertible matrices, it is enough to show that the images of $\gamma_{1},\ldots,\gamma_{N}$ generate $E^{u}$ as an $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$-module. This follows from Nakayama’s lemma. ∎

### v2

@@ 4.1. Continuity via algebra
We first focus on the case, when $f$ is a monomial $y_{1}^{d_{1}}\cdots y_{n}^{d_{n}}$ of degree $d_{1}+\dots+d_{n}=d\geq 1$. There exists an $A$-multilinear form $\beta:A^{\otimes_{A}d}\to A$, such that $f(a_{1},\dots,a_{n})=\beta(a_{1}\otimes\dots\otimes a_{1}\otimes\cdots\otimes a_{n}\otimes\cdots\otimes a_{n})$ for all $a_{1},\dots,a_{n}\in A$, where in the argument of $\beta$, $a_{i}$ is repeated $d_{i}$ times. Write $\iota:M\to A$ for the inclusion map. We have a natural $R$-linear map $\iota^{\otimes}:M^{\otimes_{R}d}\to A^{\otimes_{A}d}$ and we denote by $N$ the image of $M^{\otimes_{R}d}$ under $\tilde{\beta}:=\beta\circ\iota^{\otimes}$ in $A$. In particular $f|_{M}$ takes values in $N$ and arises as the composition
@@
Let $X=\operatorname{Spec}B$ be an affine scheme of finite type over $\mathcal{O}$. A point $x\in X(A)$ corresponds to an $\mathcal{O}$-algebra homomorphism $x:B\rightarrow A$. We thus have a canonical map
@@
    for all $n\geq 1$ and all closed immersions $\tau:X\hookrightarrow\mathbb{A}^{n}$ of $\mathcal{O}$-schemes there exists a finitely generated $R$-submodule $M\subseteq\mathbb{A}^{n}(A)=A^{n}$ such that $\tau(f(S))\subseteq M$ and $\tau\circ f:S\rightarrow M$ is continuous for the canonical $R$-module topology on the target.
@@
    there exists a closed immersion $\tau:X\hookrightarrow\mathbb{A}^{n}$ of $\mathcal{O}$-schemes and a finitely generated $R$-submodule $M\subseteq\mathbb{A}^{n}(A)=A^{n}$ such that $\tau(f(S))\subseteq M$ and $\tau\circ f:S\rightarrow M$ is continuous for the canonical $R$-module topology on the target.
@@
A closed immersion $\tau:X\hookrightarrow\mathbb{A}^{n}$ corresponds to a surjective homomorphism of $\mathcal{O}$-algebras $\mathcal{O}[x_{1},\ldots,x_{n}]\twoheadrightarrow B$. Let $b_{1},\ldots b_{n}$ be the images of $x_{1},\ldots x_{n}$ in $B$. Then the map $X(A)\overset{\tau}{\longrightarrow}\mathbb{A}^{n}(A)=A^{n}$ factors as $X(A)\rightarrow A^{B}\overset{\prod_{i=1}^{n}p_{b_{i}}}{\longrightarrow}A^{n}$. The image of $\prod_{b}M_{b}$ under this map is $\prod_{i=1}^{n}M_{b_{i}}$, which is a finitely generated $R$-module. Moreover, this projection is continuous. Hence, part (1) implies part (2).
@@
Let us assume that (3) holds. Let $M_{i}\subseteq A$ be the image of $M\subseteq\prod_{i=1}^{n}A$ under the projection onto the $i$-th component. Then $M\subseteq\prod_{i=1}^{n}M_{i}$ and $\prod_{i=1}^{n}M_{i}$ is a finitely generated $R$-module. Hence, we may replace $M$ with $\prod_{i=1}^{n}M_{i}$. Let $b_{i}\in B$ be the images of $x_{i}$ as before. If $b\in B$ then there is $g\in\mathcal{O}[x_{1},\ldots,x_{n}]$ such that $b=g(b_{1},\ldots,b_{n})$. The image of $\prod_{i=1}^{n}M_{i}$ under the map $\varphi_{g}:A^{n}\rightarrow A$, $(a_{1},\ldots,a_{n})\mapsto g(a_{1},\ldots,a_{n})$ is contained in a finitely generated $R$-module $M_{b}$ by 4.1. Since $\tau(f(S))\subseteq\prod_{i=1}^{n}M_{i}$ we have $\varphi_{g}(\tau(f(S)))\subseteq M_{b}$. Moreover, the induced map $S\rightarrow M_{b}$ is continuous since polynomial maps are continuous. The composition $\varphi_{g}\circ\tau:X(A)\rightarrow A$ is the map given by evaluation at $b$. Hence, (3) implies (1). ∎
@@ 4.2. Continuity via condensed sets
We will now express the equivalent conditions in Lemma 4.2 in terms of condensed sets. Recall that these are certain functors from the category of profinite sets to the category of sets, see [53] and Appendix A. We may view profinite sets as condensed sets via the Yoneda embedding $S\mapsto\underline{S}$, where $\underline{S}(T):=\mathcal{C}(T,S)$. The underlying set of a condensed set is by definition that functor evaluated at a point.
Recall further, that every profinite group (ring) is naturally a condensed group (ring), as the functor from topological spaces to condensed sets commutes with finite limits [53, Proposition 1.7]. We can also equip a profinite ring $R$ with the discrete topology and write $R_{\operatorname{disc}}$ to denote the associated discrete condensed ring. We have a natural map of condensed rings $R_{\operatorname{disc}}\to\underline{R}$. An abstract $R$-module $M$ can be seen as a discrete condensed $R_{\operatorname{disc}}$-module $M_{\operatorname{disc}}$, so that $M_{\operatorname{disc}}(T)=\mathcal{C}(T,M)$, where the target $M$ is given the discrete topology. The tensor product of condensed $R_{\operatorname{disc}}$-modules $M_{\operatorname{disc}}\otimes_{R_{\operatorname{disc}}}\underline{R}$ is a condensed $R$-module. [^3] Fargues–Scholze in [28, Section VIII] call such modules relatively discrete $R$-modules.
@@
We will say that a map of sets $f:S\rightarrow X(A)$ is $R$-condensed if there is a morphism of condensed sets $\tilde{f}:\underline{S}\to X(A_{\operatorname{disc}}\otimes_{R_{\operatorname{disc}}}\underline{R})$, whose map on underlying sets is $f$.
@@
By A.8 the map $X(A_{\operatorname{disc}}\otimes_{R_{\operatorname{disc}}}\underline{R})\to\mathbb{A}^{n}(A_{\operatorname{disc}}\otimes_{R_{\operatorname{disc}}}\underline{R})$ is a quasi-compact injection. It follows from [52, Proposition 4.13], that
@@ 4.3. $R$-condensed representations
(27) $\underline{\Gamma}\times\underline{\Gamma}$$G(A_{\operatorname{disc}}\otimes_{R_{\operatorname{disc}}}\underline{R})\times G(A_{\operatorname{disc}}\otimes_{R_{\operatorname{disc}}}\underline{R})$$\underline{\Gamma}$$G(A_{\operatorname{disc}}\otimes_{R_{\operatorname{disc}}}\underline{R})$$\scriptstyle{\tilde{\rho}\times\tilde{\rho}}$$\scriptstyle{\tilde{\rho}}$
@@
(1) implies (2). Let $\gamma_{1},\dots,\gamma_{N}\in\Gamma$ be generators of a dense subgroup of $\Gamma$. By assumption there is an $R$-subalgebra $B\subseteq A$, which contains an $R$-finite open $R$-subalgebra $B^{0}$, such that $\rho(\gamma_{1}),\dots,\rho(\gamma_{N})\in G(B)$. Since $G(B)$ is closed in $G(A)$, we have $\rho(\Gamma)\subseteq G(B)$. Since $G$ is of finite type over $\mathcal{O}$, $G(B^{0})$ is an open subgroup of $G(B)$. Thus the preimage $\Gamma^{\prime}:=\rho^{-1}(G(B^{0}))$ is an open subgroup of $\Gamma$. If $\tau:G\rightarrow\mathbb{A}^{n}$ is a closed immersion then $\tau(\rho(\Gamma^{\prime}))$ is contained in $\mathbb{A}^{n}(B^{0})$, which a finitely generated $R$-module. Thus the restriction of $\rho$ to $\Gamma^{\prime}$ is $R$-condensed and Lemma 4.13 implies that $\rho$ is $R$-condensed.
@@
Clearly (1) implies (2). For the converse, let $\Delta$ be the abstract subgroup of $\Gamma$ generated by $\{\gamma_{i}\}_{i\in I}$. Let $\tau:G\hookrightarrow\mathbb{A}^{n}$ be a closed immersion. Since $\rho$ is $R$-condensed $\rho(\Gamma)$ is contained in a finitely generated $R$-submodule $M\subseteq\mathbb{A}^{n}(A)$ and the map $\rho:\Gamma\rightarrow M$ is continuous. The assumption in (2) implies that $\rho(\Delta)\subseteq H(A)$. Since $H$ is closed in $G$ the restriction of $\tau$ to $H$ is a closed immersion. It follows from Lemma 4.1 that $M\cap H(A)$ is a closed subset of $M$. Its preimage in $\Gamma$ will be a closed subset of $\Gamma$ containing $\Delta$. Since $\Delta$ is dense in $\Gamma$ we conclude that $\rho(\Gamma)\subseteq M\cap H(A)$, which implies part (1). ∎
@@ 5. The space of generic matrices
Let $G$ be a smooth affine $\mathcal{O}$-group scheme and let $\overline{\rho}:\Gamma\rightarrow G(k)$ be a continuous representation, where $\Gamma$ is a profinite group satisfying Mazur’s $p$-finiteness condition. We will introduce a space of generic matrices [^4] The terminology is taken from [6] and is inspired by Procesi’s paper [49]. for $G$, which will depend on $\overline{\rho}$ and a faithful representation $\tau:G\to\operatorname{GL}_{d}$.
@@ 5.1. Definition of $X^{\mathrm{gen}}_{\operatorname{GL}_{d}}$ and $X^{\mathrm{gen},\tau}_{G}$
We fix [^5] Its existence follows from [30, Exp. $\mathrm{VII}_{B}$, Rem. 11.11.1]. a closed immersion $\tau:G\to\operatorname{GL}_{d}$ of $\mathcal{O}$-group schemes. By composing it with $\overline{\rho}$ we obtain a continuous representation $\tau\circ\overline{\rho}:\Gamma\rightarrow\operatorname{GL}_{d}(k)$. The linearisation $(\tau\circ\overline{\rho})^{\mathrm{lin}}:k[\Gamma]\to M_{d}(k)$ extends to the completed group ring $(\tau\circ\overline{\rho})^{\mathrm{lin}}:k\llbracket\Gamma\rrbracket\to M_{d}(k)$. Let $\overline{D}:k\llbracket\Gamma\rrbracket\rightarrow k$ be the continuous determinant law associated to $(\tau\circ\overline{\rho})^{\mathrm{lin}}$. (See [15] and [62, Section 3].)
Let $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$ be the universal deformation ring of $\overline{D}$ introduced in [15, Section 3.1] and let $D^{u}:R^{\mathrm{ps}}_{\operatorname{GL}_{d}}\llbracket\Gamma\rrbracket\rightarrow R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$ be the extension of the universal $d$-dimensional determinant law lifting $\overline{D}$. In particular, $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$ is a complete local noetherian $\mathcal{O}$-algebra with residue field $k$, and we equip it with its $\mathfrak{m}$-adic topology. We will write $D^{u}_{|A}$ for the specialisation of $D^{u}$ along $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}\rightarrow A$. We let $X^{\mathrm{ps}}_{\operatorname{GL}_{d}}:=\operatorname{Spec}(R^{\mathrm{ps}}_{\operatorname{GL}_{d}})$. Note, that $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$ corresponds to $R^{\mathrm{ps}}$ of [6] when we take the residual representation to be $\tau\circ\overline{\rho}$.
Let $\mathrm{CH}(D^{u})$ be the two-sided ideal of $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}\llbracket\Gamma\rrbracket$ defined in [15, Section 1.17], so that $E^{u}:=R^{\mathrm{ps}}_{\operatorname{GL}_{d}}\llbracket\Gamma\rrbracket/\mathrm{CH}(D^{u})$ is the largest quotient of $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}\llbracket\Gamma\rrbracket$ on which $D^{u}$ descends to a Cayley–Hamilton determinant law. Since $\Gamma$ satisfies Mazur’s $p$-finiteness condition, $E^{u}$ is a finitely generated $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$-module and $\mathrm{CH}(D^{u})$ is closed by [62, Proposition 3.6]. We write $D^{u}_{E^{u}}:E^{u}\to R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$ for the determinant law induced on $E^{u}$ by $D^{u}$ and $\pi_{E^{u}}:R^{\mathrm{ps}}_{\operatorname{GL}_{d}}\llbracket\Gamma\rrbracket\to E^{u}$ for the projection.
@@
It follows from [62, Proposition 3.6] that $E^{u}$ is a finitely generated $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$-module. This implies that the quotient topology on $E^{u}$ coincides with the topology inherited from $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$ as a finitely generated $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$-module, see the proof of [6, Lemma 3.2]. Every $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$-submodule of $E^{u}$ is closed with respect to this topology. Hence, $S$ is closed in $E^{u}$, and thus $S=E^{u}$. ∎
@@
(1) implies (2). The homomorphism $\widehat{\rho}\otimes A:R^{\mathrm{ps}}_{\operatorname{GL}_{d}}\llbracket\Gamma\rrbracket\otimes_{R^{\mathrm{ps}}_{\operatorname{GL}_{d}}}A\to M_{d}(A)$ induces a $d$-dimensional AA-valued Cayley–Hamilton determinant law
@@
Part (1). Recall, that in [6, Lemma 3.1] the $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$-algebra $A^{\mathrm{gen}}$ is proved to represent the functor, that maps an $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$-algebra $A$ to the set of homomorphisms of Cayley–Hamilton algebras $f:E^{u}\to M_{d}(A)$. Once we prove this functor to be naturally isomorphic to $X^{\mathrm{gen}}_{\operatorname{GL}_{d}}$ as in 5.3, the claim about representability follows and $X^{\mathrm{gen}}_{\operatorname{GL}_{d}}$ is naturally isomorphic to the $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$-scheme $X^{\mathrm{gen}}=\operatorname{Spec}A^{\mathrm{gen}}$ from [6].
If $\rho\in X^{\mathrm{gen}}_{\operatorname{GL}_{d}}(A)$ then by Lemma 5.2 (2) we obtain a natural homomorphism of Cayley–Hamilton algebras $f_{\rho}:E^{u}\to M_{d}(A)$. Conversely, a map of Cayley–Hamilton algebras $f:E^{u}\to M_{d}(A)$ gives rise to a representation by means of the composition $\rho_{f}:\Gamma\to(E^{u})^{\times}\overset{f}{\to}\operatorname{GL}_{d}(A).$ By restriction along $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}\llbracket\Gamma\rrbracket\otimes_{R^{\mathrm{ps}}_{\operatorname{GL}_{d}}}A\to E^{u}\otimes_{R^{\mathrm{ps}}_{\operatorname{GL}_{d}}}A$, we see that
@@
Part (2) follows from [6, Lemma 3.2] or Lemma 4.9. ∎
@@
As already explained above, [62, Proposition 3.6] implies that $E^{u}$ is a finitely generated $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$-module, and hence $k\otimes_{R^{\mathrm{ps}}_{\operatorname{GL}_{d}}}E^{u}$ is a finite dimensional $k$-vector space.
@@
It follows from Part (1) of Lemma 5.4 that $X^{\mathrm{gen}}_{\operatorname{GL}_{d}}$ coincides with the scheme $X^{\mathrm{gen}}$, that was contructed in [6, Lemma 3.1] as a closed subscheme of $M_{d}^{n}\times X^{\mathrm{ps}}_{\operatorname{GL}_{d}}$ by choosing generators $\delta_{1},\ldots,\delta_{n}$ of $E^{u}$ as an $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$-module and mapping $\rho$ to the $n$-tuple $(\rho(\delta_{1}),\ldots,\rho(\delta_{n}))$. Since $\rho(\gamma_{i})$ are invertible matrices, it is enough to show that the images of $\gamma_{1},\ldots,\gamma_{N}$ generate $E^{u}$ as an $R^{\mathrm{ps}}_{\operatorname{GL}_{d}}$-module. This follows from Nakayama’s lemma. ∎